OpenStudy (anonymous):

When you're just adding coefficants with exponants what do you add? For example: 7x^2+9x-3x^3

OpenStudy (anonymous):

If you add terms with the same variable and exponent, then you add just the coefficients: 3x^2+5x^2=(3+5)x^2=8x^2. When you multiply like bases (same variable) then you add exponents: x^2 * x^3 = x^(2+3)=x^5.

OpenStudy (anonymous):

So right now I'm just adding Coefficients with random and different exponents. How would I solve this?

OpenStudy (anonymous):

Give me another example. Your original example is already simplified because there are no "like" or "similar" terms; neither are there any multiplications.

OpenStudy (anonymous):

Well In this question they just asked me to simplify. So i guess no work needs to be done then?

OpenStudy (anonymous):

Sometimes with a polynomial, the book or your teacher wants the exponents to be decreasing in power. So you would then write 7x^2+9x-3x^3 as -3x^3+7x^2+9x. Otherwise, there is no other simplification possible.

OpenStudy (anonymous):

thanks.